When CBI Corporation priced notebooks at P50 each, it

sold 35 notebooks per month. Demand increased for notebooks raising the price

of it to P55 each. CBI responded by increasing its quantity supplied to 42 notebooks

per month. Using the point method for calculating elasticity, what is CBI’s

price elasticity of supply?




1
Expert's answer
2021-09-10T12:21:39-0400
"\frac{p-50}{55-50}=\frac{Q-35}{42-35}"

"Q=\frac{7}{5}p-\frac{175}{5}"

"E=\frac{7}{5}\times\frac{55}{42}=1\frac{5}{6}"